An object carries a charge q=-4.8*10 -6 C. calculate the number of electrons corresponding to this charge
1
Expert's answer
2020-10-15T10:50:06-0400

The charge on a single electron is "q_e=1.6\times 10^{−19}\;C" . All other charges in the universe consist of an interger multiple of this charge ("q=nq_e" ). This is known as charge quantisation.

A negatively charged object has an excess of electrons on its surface.

"n = \frac{q}{q_e}"

"n = \frac{-4.8 \times 10^{-6}}{1.6\times 10^{-19}}"

"n = 3\times 10^{13}"

Answer: 3×1013